OpenStudy (anonymous):

sofia made 35 percent of her shots in a basketball game.what is the ratio of the number of shots sofia made to the number of shots she did NOT make in the game?

hero (hero):

Solution: Sofia made 35% of her shots, therefore she missed 65% of her shots. Since 35% = 35/100 and 65% = 65/100, the ratio of the number of shots made to the number of shots missed is 35:65 or 7:13 or 7/13 in fraction form.

hero (hero):

Any questions?

OpenStudy (anonymous):

Nope I Understand thanks!